Using two even consecutive integers where two times the lower integer plus the higher integer equals 50.
I have tried: 2x2+4=8, 4x4+6=22, 6x6+8=44, 8x8+10=64 so I can get close but not equal to 50 without going over

Answer by nerdybill(7384)

Using two even consecutive integers where two times the lower integer plus the higher integer equals 50.
.
Let x = lower even integer
then
x+2 = higher consecutive even integer
.
2x + (x+2) = 50
3x + 2 = 50
3x = 48
x = 16 (lower)
.
higher:
x+2 = 16+2 = 18
.
answer: 16 and 18